# suppSupplement

### v1

@@ E.8 Noise analysis
$\langle\hat{a}^{{\dagger},\mathrm{in}}_{j}(t)\hat{a}^{\mathrm{in}}_{j}(t^{\prime})\rangle$ $=n_{j}^{\mathrm{in}}\delta(t-t^{\prime})$ (43a)
$\langle\hat{a}^{\mathrm{in}}_{j}(t)\hat{a}^{{\dagger},\mathrm{in}}_{j}(t^{\prime})\rangle$ $=(n_{j}^{\mathrm{in}}+1)\delta(t-t^{\prime})$ (43b)
$\langle\hat{a}^{{\dagger},\mathrm{noise}}_{j}(t)\hat{a}^{\mathrm{noise}}_{j}(t^{\prime})\rangle$ $=n_{j}^{\mathrm{noise}}\delta(t-t^{\prime})$ (43c)
$\langle\hat{a}^{\mathrm{noise}}_{j}(t)\hat{a}^{{\dagger},\mathrm{noise}}_{j}(t^{\prime})\rangle$ $=(n_{j}^{\mathrm{noise}}+1)\delta(t-t^{\prime})$ (43d)
@@
with $n_{j}^{\mathrm{in}}$ and $n_{j}^{\mathrm{noise}}$ as the occupation number of the baths connected to the respective mode.
@@
The number of added noise photons on the input (mode 1) and output channel (mode 2) equals:
@@
$N_{\mathrm{input}}^{\mathrm{add}}=\frac{1}{2}\sum_{k}\left((n_{k}^{\mathrm{in}}+1)\absolutevalue{S_{1k}}^{2}+(n_{k}^{\mathrm{noise}}+1)\absolutevalue{N_{1k}}^{2}\right)$ (44a)
@@
$N_{\mathrm{output}}^{\mathrm{add}}=\frac{1}{2}\sum_{k\neq 1}\left((n_{k}^{\mathrm{in}}+1)\absolutevalue{S_{2k}}^{2}+(n_{k}^{\mathrm{noise}}+1)\absolutevalue{N_{2k}}^{2}\right)$ (44b)
@@
The quantum limit for a phase-preserving amplifiers is $N_{\mathrm{input}}^{\mathrm{add}}=1/2$ and $N_{\mathrm{output}}^{\mathrm{add}}=(G-1)/2$ [39]. We assume that all the baths are at zero temperature, except the bath connected to the output channel. So, $n_{j}^{\mathrm{in}}=0$ if $j\neq 2$, $n_{j}^{\mathrm{noise}}=0$ for all $j$, and $n_{2}^{\mathrm{in}}>0$.
During optimization, we set $N_{\mathrm{input}}^{\mathrm{add}}=1/2$ as an additional constraint during the optimization. Therefore, all identified architectures fulfill this equality constraint. The quantum limit for the output channel was not as additional constraint and is analysed in the following.
For Graph (I,II,UI) the number of added photons on the output channel turn out to equal $N_{\mathrm{output}}^{\mathrm{add}}=(G-1)/2$, therefore, reaching the quantum limit for all $G$. For the graphs (III–V,UII) $N_{\mathrm{output}}^{\mathrm{add}}=(G+1)/2$. So, the number of added photons is close to the quantum limit and approaches it in the limit of $G\to\infty$.
Appendix F Bus modes
Here, we give more details on how we have augmented our standard method to treat the bus modes in optomechanical and other hybrid systems that feature constraints in the connectivity of the port modes. These modifications are introduced to facilitate the discovery of schemes that require infinite cooperativity between the bus modes and other modes to reach the ideal behavior.
F.1 Far-detuned bus modes
To discover schemes in which one or more bus modes mediate purely coherent interactions between the port modes we introduce as an additional building block of our discrete optimization a far-detuned bus mode. Formally, we set the decay rates of this type of mode to zero and at the same time assume that their detuning is not zero. This allows us to eliminate all far-detuned modes and incorporate their effect into an effective Hamiltonian for the remaining modes, as shown below.
For $N$ standard modes $\hat{a}_{j}$ and $M$ far-detuned modes $\hat{b}_{j}$ the most general quadratic Hamiltonian reads:
$\hat{H}=\frac{1}{2}\sum_{jk=1}^{NN}\left(g^{(aa)}_{jk}\hat{a}_{j}^{\dagger}\hat{a}_{k}+\nu^{(aa)}_{jk}\hat{a}_{j}^{\dagger}\hat{a}_{k}^{\dagger}\right)+\frac{1}{2}\sum_{jk=1}^{MM}\left(g^{(bb)}_{jk}\hat{b}_{j}^{\dagger}\hat{b}_{k}+\nu^{(bb)}_{jk}\hat{b}_{j}^{\dagger}\hat{b}_{k}^{\dagger}\right)+\sum_{jk=1}^{NM}\left(g^{(ab)}_{jk}\hat{a}_{j}^{\dagger}\hat{b}_{k}+\nu^{(ab)}_{jk}\hat{a}_{j}^{\dagger}\hat{b}_{k}^{\dagger}\right)+\mathrm{H.c.},$ (45)
where $g^{(aa)}$ and $g^{(bb)}$ are hermitian matrices, $\nu^{(aa)}$ and $\nu^{(bb)}$ are symmetric matrices, and $g^{(ab)}$ and $\nu^{(ab)}$ are arbitrary matrices. To compactly rewriting the effective Hamiltonian, it is convenient to define the matrices
$H^{(ss^{\prime})}_{\rm BdG}=\begin{pmatrix}g^{(ss^{\prime})}&\nu^{(ss^{\prime})}\\ \nu^{(ss^{\prime})*}&g^{(ss^{\prime})*}\\ \end{pmatrix}$ (46)
with $s,s^{\prime}=a$ or $b$ and $g^{(ba)}=g^{(ab)*}$ and $\nu^{(ba)}=\nu^{(ab)}$. These matrices can be viewed as blocks of the underlying Bogoliubov de Gennes Hamiltonian. The effective Bogoliubov de Gennes Hamiltonian can be written as
$H^{\mathrm{({\rm eff})}}_{{\rm BdG}}=H^{\mathrm{(aa)}}_{{\rm BdG}}-H^{\mathrm{(ab)}}_{{\rm BdG}}(H^{\mathrm{(bb)}}_{{\rm BdG}})^{-1}H^{\mathrm{(ba)}}_{{\rm BdG}}.$ (47)
The corresponding second-quantized Hamiltonian reads
$\hat{H}=\frac{1}{2}\sum_{jk=1}^{NN}\left(g^{(\rm eff)}_{jk}\hat{a}_{j}^{\dagger}\hat{a}_{k}+\nu^{(\rm eff)}_{jk}\hat{a}_{j}^{\dagger}\hat{a}_{k}^{\dagger}\right)+\mathrm{H.c.}$ (48)
with
$H^{({\rm eff})}_{\rm BdG}=\begin{pmatrix}g^{({\rm eff})}&\nu^{({\rm eff})}\\ \nu^{({\rm eff})*}&g^{({\rm eff})*}\\ \end{pmatrix}$ (49)
Next, we assume that the Bogoliubov de Gennes Hamiltonian $H^{(bb)}_{\rm BdG}$ projected on the far-detuned modes had been diagonalized to eliminate any coupling between the bus modes. In other words, $\nu^{(bb)}$ is zero and $g^{(bb)}$ is diagonal with $g^{(bb)}_{ii}=-\Delta^{(b)}_{i}$. In this case, we find
$g_{ij}^{(\rm eff)}$ $=$ $g_{ij}^{(aa)}+\sum_{k}\frac{g^{(ab)}_{ik}g^{(ab)*}_{kj}}{\Delta^{(b)}_{k}}+\frac{\nu^{(ab)}_{ik}\nu^{(ab)*}_{kj}}{\Delta^{(b)}_{k}}$ (50)
$\nu_{ij}^{(\rm eff)}$ $=$ $\nu_{ij}^{(aa)}+\sum_{k}\frac{g^{(ab)}_{ik}\nu^{(ab)}_{kj}}{\Delta^{(b)}_{k}}+\frac{\nu^{(ab)}_{ik}g^{(ab)}_{kj}}{\Delta^{(b)}_{k}}.$ (51)
Thus, the rescaled Bogoliubov Hamiltonian in Eq. 11a is modified as following
$H_{ij}$ $=$ $\frac{g_{ij}^{(aa)}}{\sqrt{\kappa_{i}\kappa_{j}}}+\sum_{k}{\rm sgn}(\Delta_{k})\left(\frac{g^{(ab)}_{ik}}{\sqrt{\kappa_{i}\absolutevalue{\Delta^{(b)}_{k}}}}\frac{g^{(ab)*}_{kj}}{\sqrt{\kappa_{j}\absolutevalue{\Delta^{(b)}_{k}}}}+\frac{\nu^{(ab)}_{ik}}{\sqrt{\kappa_{i}\absolutevalue{\Delta^{(b)}_{k}}}}\frac{\nu^{(ab)*}_{kj}}{\sqrt{\kappa_{j}\absolutevalue{\Delta^{(b)}_{k}}}}\right).$ (52)
or
$H_{ij}$ $=$ $\frac{\nu_{ij}^{(aa)}}{\sqrt{\kappa_{i}\kappa_{j}}}+\sum_{k}{\rm sgn}(\Delta_{k})\left(\frac{g^{(ab)}_{ik}}{\sqrt{\kappa_{i}\absolutevalue{\Delta^{(b)}_{k}}}}\frac{\nu^{(ab)}_{kj}}{\sqrt{\kappa_{j}\absolutevalue{\Delta^{(b)}_{k}}}}+\frac{\nu^{(ab)}_{ik}}{\sqrt{\kappa_{i}\absolutevalue{\Delta^{(b)}_{k}}}}\frac{g^{(ab)}_{kj}}{\sqrt{\kappa_{j}\absolutevalue{\Delta^{(b)}_{k}}}}\right).$ (53)
We use the dimensionless parameters and $g^{(ab)}_{ik}/\sqrt{\kappa_{i}\Delta^{(b)}_{k}}$ with $k=1,...,M$ and $i=1,...,N$ as additional learning parameters of our continuous optimization. The signs ${\rm sgn}(\Delta_{k})$ remain fixed during the continuous optimization and are, instead, chosen within the discrete optimization. Note the different scaling of the learning parameters encoding the interaction with a bus mode with the detuning $\Delta^{(b)}_{k}$ playing the role of the decay rate $\kappa^{(b)}_{k}$. This scaling reflects that the coherent interactions mediated by the bus modes become independent of the bus mode decay $\kappa^{(b)}_{k}$ in the asymptotic limit of large cooperativities, $4|g^{(ab)}_{ik}|^{2}/|\kappa_{i}\kappa^{(b)}_{k}|\to\infty$. Similar to the decay rates $\kappa_{i}$ of the standard modes, we can view the detunings $\Delta_{i}^{(b)}$ of the far-detuned bus mode as free parameters (possibly subject to some stability constraints).

### v2

@@ Appendix H Noise analysis
$\langle\hat{a}^{{\dagger},s}_{j}(t)\hat{a}^{s}_{j}(t^{\prime})\rangle$ $=n_{j}^{s}\delta(t-t^{\prime})$ (17a)
$\langle\hat{a}^{s}_{j}(t)\hat{a}^{{\dagger},s}_{j}(t^{\prime})\rangle$ $=(n_{j}^{s}+1)\delta(t-t^{\prime})$ (17b)
@@
with $s$ standing either for in or noise. $n_{j}^{\mathrm{in}}$ and $n_{j}^{\mathrm{noise}}$ are the occupation numbers of the baths connected to the respective mode.
@@
For the phase-preserving amplifier discussed in Fig. 4(b) the number of added noise photons on the input (mode 1) and output channel (mode 2) equals:
@@
$N_{\mathrm{input}}^{\mathrm{add}}=\frac{1}{2}\sum_{k}\left((n_{k}^{\mathrm{in}}+1)\absolutevalue{S_{1k}}^{2}+(n_{k}^{\mathrm{noise}}+1)\absolutevalue{\mathcal{N}_{1k}}^{2}\right)$ (18a)
@@
$N_{\mathrm{output}}^{\mathrm{add}}=\frac{1}{2}\sum_{k\neq 1}\left((n_{k}^{\mathrm{in}}+1)\absolutevalue{S_{2k}}^{2}+(n_{k}^{\mathrm{noise}}+1)\absolutevalue{\mathcal{N}_{2k}}^{2}\right)$ (18b)
The quantum limit for a phase-preserving amplifiers is $N_{\mathrm{input}}^{\mathrm{add}}=1/2$ and $N_{\mathrm{output}}^{\mathrm{add}}=(G-1)/2$ [40]. If the two modes are in different subsets $M_{1}$ and $M_{2}$ (see Appendix F), the added noise on the output is increased by one photon, meaning that the minimum number of added photons equals $N_{\mathrm{output}}^{\mathrm{add}}=(G+1)/2$ in this case.
@@
We assume that all the baths are at zero temperature, except the bath connected to the output channel. So, $n_{j}^{\mathrm{in}}=0$ if $j\neq 2$, $n_{j}^{\mathrm{noise}}=0$ for all $j$, and $n_{2}^{\mathrm{in}}>0$.
During optimization, we enforce that the noise on the input and output channel equals their respective quantum limits by adding the two additional terms $|f_{1}|^{2}=(N_{\mathrm{input}}^{\mathrm{add}}-1/2)^{2}$ and $|f_{2}|^{2}=(N_{\mathrm{output}}^{\mathrm{add}}-(G\pm 1)/2)^{2}$ to the cost function Eq. (3). Therefore, all identified architectures automatically fulfill these equality constraints. The sign in constraint $f_{2}$ is chosen according to the tested graph. If input and output modes are in the same subset $M_{1}$ or $M_{2}$, we choose the minus sign. Otherwise, we choose the plus sign.
Appendix I Minimal number of parametric interactions
Here, we summarize our algorithm for finding the minimum number of parametric pumps needed to realize a given graph. The key problem is to sort the graph’s nodes into frequency classes. If two modes are in the same class, they have the same frequency, otherwise, they do not. Each graph edge connecting modes of two different classes requires a pump. The goal is to minimize the number of such edges under the constraint that the modes coupled by squeezing (blue) or complex beamsplitter interactions (green) must be in different classes.
We solve this problem in the following way: Iterating through all nodes, we assign to the current node either one of the frequency classes already assigned to a previous node, or we assign a new one, which was not used before. After all nodes have been labeled, we test whether modes coupled by squeezing (blue) or complex beamsplitter interactions (green) are in different classes. If this is true, we count the number of pumps. Otherwise, this labeling is skipped. We repeat this algorithm until every possible distribution of frequency classes has been tested. This search yields the minimum number of pumps, and also all possible combinations of carrier frequencies that must be chosen equally to each other so that this minimum is reached.
The task described above is in fact a so-called “graph coloring” problem. In our setting, the frequency class plays the role of the “color”.
Appendix J Gauge freedom and Gauge fixing
The scattering matrix is a complex matrix that encodes the amplitudes and phases of the output fields in response to a set of input fields. Many experiments focus on the amplitude of the fields. Nevertheless, the phases can also be observed after fixing a gauge by committing to specific experimental settings and physical interpretations of the fields $\hat{a}_{j}^{\rm in/out}$. However, we do not want to distinguish between phases that depend on the detailed arrangement of the experimental setup or equivalent physical interpretations. For this purpose, we allow for the free parameters $\gamma^{\rm in/out}_{j}$ in our cost function Eq. (3). The goal of this appendix is to make an explicit connection between this gauge freedom and the aforementioned experimental settings and physical interpretations.
We start discussing a gauge freedom that depends on the detailed arrangement of the experimental setup. These are the reference positions where the input and output fields are evaluated. Assume that we change the reference position for port $j$ and that the new reference position is further away from the device. An incoming signal acquires the additional phase $\phi_{j}$ while traveling from the new reference point to the old reference point. The same phase will also be acquired by an outgoing signal when traveling from the old to the new reference point. In other words, the transfer matrix will change according to the rule $S_{jk}\to S_{jk}e^{i(\phi_{j}+\phi_{k})}$ where $\phi_{j}+\phi_{k}$ is the overall additional phase acquired by a signal traveling from the new reference position in port $k$ to the new reference position in port $j$. Comparing to Eq. (3), we see that $\gamma_{j}^{\rm out}=-\gamma_{j}^{\rm in}=\phi_{j}$.
An additional gauge freedom for the fields $\hat{a}_{j}^{\rm in/out}$ comes from the arbitrary phase $\phi_{j}$ in the transformation to the rotating frame, $\hat{a}^{\rm in/out}_{j}=\exp[i(\omega_{L,j}t+\phi_{j})]\hat{a}^{\rm in/out}_{{\rm lab},j}$ where $\hat{a}^{\rm in/out}_{{\rm lab},j}$ are the fields in the lab frame. The phase $\phi_{j}$ determines the physical interpretation of the quadratures $\hat{X}^{\rm in/out}_{j}$ and $\hat{P}^{\rm in/out}_{j}$ (real and imaginary part of the field),
$\hat{X}_{j}^{\rm in}=\hat{X}^{\rm in}_{{\rm lab},j}\cos(\omega_{L,j}t+\phi_{j})-\hat{P}^{\rm in}_{{\rm lab},j}\sin(\omega_{L,j}t+\phi_{j})$ (19)
The quadratures of the field in the lab frame have the same physical interpretation at all ports, e.g., in electromagnetism $\hat{X}^{\rm in}_{{\rm lab},j}$ could correspond to the amplitude of an electric field with a fixed polarization and $\hat{P}^{\rm in}_{{\rm lab},j}$ to the amplitude of the transverse magnetic field. For incommensurate carrier frequencies $\omega_{L,j}$, any special initial choice of phases $\phi_{j}$, e.g., $\phi_{j}=\phi$, would quickly dephase. As a consequence, all gauges $\phi_{j}$ are physically equivalent. To set them on equal footing we allow for arbitrary $\gamma_{j}^{\rm out}=\gamma_{j}^{\rm in}=\phi_{j}$ in Eq. 3. Combining this gauge freedom with the choice of a reference position in each port waveguide we arrive at independent values for $\gamma_{j}^{\rm out}$ and $\gamma_{j}^{\rm in}$ in Eq. 3.
In the special case of equal carrier frequency $\omega_{L,j}=\omega_{L}$, the uniform gauge $\phi_{j}=\phi$ has a special physical meaning. In this gauge, the physical interpretation of the quadratures depends on time but remains the same at every port, see Eq. 19. In this experimental scenario, the target scattering matrix is usually written in the uniform gauge and one should restrict the gauge freedom to $\gamma_{j}^{\rm out}=-\gamma_{j}^{\rm in}=\phi_{j}$. To highlight the importance of the right choice for the gauge freedom we consider a concrete example in the gyrator, see Supplemental Material G. This is a non-reciprocal device with scattering matrix
$S_{\mathrm{target}}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}.$ (20)
For a gyrator, it is assumed that the carrier frequencies and the physical meaning of the fields at the two ports are the same. With this assumption, we can interpret the scattering matrix as describing a non-reciprocal device in which a field injected into the first port exits the second port as the same field, e.g., an electric field, but with an opposite sign. On the other hand, exchanging the emitter and the receiver, a field entering from the second port exits the first port without any sign change.
On the contrary, without the assumption that the quadratures have the same physical interpretation at both ports, the sign difference between the two off-diagonal elements loses any meaning. Indeed, it can be gauged away with the appropriate choice of $\gamma_{j}^{\rm out}$ and $\gamma_{j}^{\rm in}$ in Eq. (3), e.g. $\gamma_{2}^{\rm out}=\pi$ and $\gamma_{2}^{\rm in}=\gamma_{1}^{\rm in/out}=0$.
As the choice of gauge freedom is platform dependent, AutoScatter allows the user to choose between four different options for handling the gauge freedoms: (i) no gauge freedom, (ii) allow only for free reference positions of the ports ($\gamma_{j}^{\rm out}=-\gamma_{j}^{\rm in}$), (iii) allow only for the gauge freedoms of the fields $\hat{a}_{j}^{\mathrm{in/out}}$ ($\gamma_{j}^{\rm out}=\gamma_{j}^{\rm in}$), and (iv) allow for both degrees of freedom (independent $\gamma_{j}^{\rm out}$ and $\gamma_{j}^{\rm in}$).
In all examples presented in the main text, one could use both options (ii) and (iv) to arrive at the same results. However, as we discussed using the gyrator as an example, option (ii) is sometimes implicitly assumed for certain non-reciprocal devices. For this reason, we have used this option to produce the data in Fig. 7 and Supplemental Material F and G.